# SeaFormer++: Squeeze-enhanced Axial Transformer for Mobile Visual Recognition

### v5

@@ 8.3 Ablation studie
Cls loss Out loss Feat loss Cross loss mIoU
32.5
33.7
34.7
35.5
Table 16: Ablation study results of loss function design.
Through ablation studies, this experiment evaluates the performance impact of different loss functions in semantic segmentation, involving ”classification loss,” ”output similarity loss,” ”feature similarity loss,” and ”cross-model classification loss,” with mIoU as the evaluation metric. Starting from a baseline mIoU of 32.5 with just classification loss, performance sequentially improves with the addition of output and feature similarity losses, highlighting the benefits of aligning student and teacher model outputs and features for accuracy. The highest mIoU of 35.5 is achieved with the inclusion of cross-model classification loss, emphasizing the effectiveness of combining various constraints on the student model. The results underscore the individual and collective contributions of each loss function to semantic segmentation tasks, particularly the significance of model alignment for substantial performance gains, and the synergistic enhancement of model performance through integrated loss function design.
@@ 10 Data Availability Statement
The datasets generated during and/or analysed during the current study are available in the Imagenet [63] (https://www.image-net.org/), COCO [61] (https://cocodataset.org), ADE20K [58] (https://groups.csail.mit.edu/vision/datasets/ADE20K/), Cityscapes [59] (https://www.cityscapes-dataset.com), Pascal Context [60] (https://cs.stanford.edu/~roozbeh/pascal-context/) and COCO-Stuff [61] (https://github.com/nightrome/cocostuff?tab=readme-ov-file) repositories.

### v6

@@ 8.3 Ablation studies
Through ablation studies, this experiment evaluates the performance impact of different loss functions in semantic segmentation, involving classification loss, output similarity loss, feature similarity loss, and cross-model classification loss, with mIoU as the evaluation metric. Starting from a baseline mIoU of 32.5 with just classification loss, performance sequentially improves with the addition of output and feature similarity losses, highlighting the benefits of aligning student and teacher model outputs and features for accuracy. The highest mIoU of 35.5 is achieved with the inclusion of cross-model classification loss, emphasizing the effectiveness of combining various constraints on the student model. The results underscore the individual and collective contributions of each loss function to semantic segmentation tasks, particularly the significance of model alignment for substantial performance gains, and the synergistic enhancement of model performance through integrated loss function design. To demonstrate the necessity and effectiveness of multi-resolution distillation. We compare with the conventional distillation baseline with a low-resolution teacher. As seen in the table, mIoU of the low-resolution teacher method is 32.1, which is significantly lower than the 35.5 achieved by our multi-resolution distillation method. This highlights the necessity and effectiveness of multi-resolution distillation for improving model performance.
@@ 10 Data Availability Statement
The datasets generated during and/or analysed during the current study are available in the Imagenet Deng et al. (2009) (https://www.image-net.org/), COCO Caesar et al. (2018) (https://cocodataset.org), ADE20K Zhou et al. (2017) (https://groups.csail.mit.edu/vision/datasets/ADE20K/), Cityscapes Cordts et al. (2016) (https://www.cityscapes-dataset.com), Pascal Context Mottaghi et al. (2014) (https://cs.stanford.edu/~roozbeh/pascal-context/) and COCO-Stuff Caesar et al. (2018) (https://github.com/nightrome/cocostuff?tab=readme-ov-file) repositories.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).